OpenStudy (anonymous):

What is the slope of the line passing through the points (0, −5) and (4, 2)?

OpenStudy (solomonzelman):

\(\LARGE \color{black}{ \displaystyle {\rm m}=\frac{\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2}}{\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2}} }\) where: \(\LARGE \color{black}{ \displaystyle {\rm m} }\) is the slope \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (anonymous):

Right I get that and I get the formula

OpenStudy (alexandervonhumboldt2):

substitute the coordinates in the formula soloman gave you and solve

OpenStudy (alexandervonhumboldt2):

(-5-2)/(0-4)=solve

OpenStudy (anonymous):

When I did that I got -7 and -4 so would it be -7/-4?